OpenStudy (anonymous):

draw a number line to find each sum (1) -3+2 (2) -5+4 (3) 8+(-2)

OpenStudy (anonymous):

help please

OpenStudy (anonymous):

-3+2=-1 -5+4=-1 8+(-2)=6